5x+12-4x-14=22
9x+12-14=22
9x+-2=22
+2=+2
9x=24
9x/3=24/3
x=3 but I'm thinking this is wrong?

Answer by Alan3354(69443)

5x+12-4x-14=22
9x+12-14=22
9x+-2=22
+2=+2
9x=24 ************** OK to here
---------
9x/3=24/3 This is not wrong
--> 3x = 8
3x/3 = 8/3
x = 8/3
===============
However:
5x+12-4x-14=22
9x+12-14=22 ********* not 9x
x+12-14 = 22
x - 2 = 22
x = 24